Use a graph of the Keynesian cross to show the effects of an increase in exogenous planned investment on the equilibrium level of income/output. Be sure to label: i. the axes; ii. the curves; iii. the initial equilibrium values; iv. the direction the curve shifts; and v. the terminal equilibrium values.
Answer #1

The planned expenditure (in a closed economy) is PE = C + I + G , and for the consumption function be C = Co - mpc(Y-T) , we have PE = Co + mpc(Y - T) +I+G or PE = (Co- mpc.T+I+G) + mpc.Y , where mpc is marginal propensity to consume, which is between 0 and 1. Also, mpc would be the slope while the intercept would be Co- mpc. T + 1 + G) . For an increase in planned investment 'I', the new planned expenditure would be PE = (Co- mpc.T+1 +G) + mpc.Y . The slope does not change, but the intercept increases by I'-I, which is the amount by which the PE curve shifts left.

The graph would be as below.

1000 PE 1800- 1600- New +400-PE Old PE 2004 PE=Y Y 1000 200 400 600 800

As can be seen, output is on horizontal axis while PE is on vertical axis. The initial PE is taken as PE = 200 + 0.6Y , and for an increase in investment by 100, we have the new PE as PE = 300 + 0.6Y . The equilibrium (where Y=PE) shifts from E to E', for the shift in the PE curve (red line) upward. Also, the blue line presents the 45 degree line of PE=Y.